OpenStudy (anonymous):

[(3800x-(80698/967432*3.14))^(1/3)/(3.14x-88x+(400/8009)^2]*[((4/90)(8x/32)^(1/4)*88^(2)*x^(3)-32^(8))/((90/4)(32/x)^(3)-x+x^(8)-(55/8)]-(45x/88e^(10))^(1/2)=0

OpenStudy (anonymous):

your paranthesis do not match. use the equation editor below to show what it is!

OpenStudy (anonymous):

now, the first part of it is chopped off

OpenStudy (anonymous):

and how on earth did you end up with that gianta**question?

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{3800x-\frac{ 80698 }{ 967432\Pi }} }{ (\Pi*x-88x(\frac{ 400 }{ 8009 }))^{2} } *[\frac{ \frac{ 4 }{ 90 }\ \sqrt[4]\frac{ 8x }{ 32 }*88^{2}*x ^{3}-32^{8} }{\frac{ 90 }{ 4 }*(\frac{ 32 }{ x })^{3}-x+x ^{8}-\frac{ 55 }{ 8 } }=\sqrt{\frac{ 45x }{ 88e ^{10} }}\]

OpenStudy (anonymous):

@electrokid wanna come help??

OpenStudy (anonymous):

wanna help @hea?

OpenStudy (anonymous):

I would smash that into wolfram alpha to be honest, it's a little too big to do in my head. Sorry.